OpenStudy (anonymous):

Find the slope of the tangent line to the curve (a lemniscate) 2(x^2+y^2)^2=25(x^2-y^2) at the point (-3,1)?

hartnn (hartnn):

could you implicitly differentiate that ?

hartnn (hartnn):

you will need chain rule here whats d/dx x^2 =...? whats d/dx y^2=...?

hartnn (hartnn):

what ? where does 2*2 come from ?

hartnn (hartnn):

d/dx[2(x^2+y^2)^2] = 2*(2 (x^2+y^2)) times d/dx(x^2+y^2) this is because of chain rule, got this step ?

hartnn (hartnn):

d/dx (x^2) = 2x thats why and the other 2 was already there

hartnn (hartnn):

ok, so now what about d/dx (x^2+y^2) =...?

OpenStudy (anonymous):

2x+2y?

hartnn (hartnn):

here, chain rule again comes into picture d/dx f(y) = f'(y) dy/dx so d/dx (y^2) = 2y dy/dx got this ?

hartnn (hartnn):

if you got that we are almost done :)

hartnn (hartnn):

d/dx[2(x^2+y^2)^2] = 2*(2 (x^2+y^2)) times d/dx(x^2+y^2) this part ?

hartnn (hartnn):

whats d/dx x^2 =.... ?

hartnn (hartnn):

we don't rewrite anything

OpenStudy (anonymous):

ohman :/

hartnn (hartnn):

derivative of outer times derivative of inner like for (sin x)^3 we have outer = x^3, derivative = 3x^2 inner = sin x derivative = cos x so, 3 (sin x)^2 cos x got this ?

hartnn (hartnn):

in same way d/dx [(x^2+y^2)^2] = 2 (x^2+y^2) times d/dx (x^2+y^2)

OpenStudy (anonymous):

i get that, but in this problem im confused with why the exponent goes away

hartnn (hartnn):

ok, you said d/dx (x^2) = 2x why did the exponent go away here ?

OpenStudy (anonymous):

2(x^2+y^2)^2=25(x^2-y^2)

OpenStudy (anonymous):

so you're finding the derivative of the inside?

hartnn (hartnn):

we do find derivative of inside in chain rule, right ?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

my professor also taught us a way of outer first then rewrite inner and then multiple by the derivative of the inner.

hartnn (hartnn):

he must have meant "derivative of outer first ....."

OpenStudy (anonymous):

d/dx[f(g(x))]= f' (g(x))*g'(x)

hartnn (hartnn):

yes, here f(x) was x2

OpenStudy (anonymous):

well isn't it 4?

OpenStudy (anonymous):

2(x^2+y^2)^2